# Far from Perfect: Quantum Error Correction with (Hyperinvariant) Evenbly Codes

### v1

@@ IV.2 Error Correction
The results conveyed demonstrate that Evenbly codes achieve similar logical recovery rates when compared to both topological and quantum low-density parity-check (qLDPC) codes [7, 6] in the code-capacity regime, as well as other holographic codes [13, 22, 28, 9, 8]. However, in contrast to the surface or color codes [41], the distance of all bulk logical qubits scale as $\sim n^{2/3}$, rather than $\sim\sqrt{n}$; what is more, the rate of the code is tunable, a property generally ascribed to holographic codes [8, 14]. It would be interesting to examine phenomenological noise models (which incorporate measurement errors); we leave this idea for future work.
@@
Additionally, we reported that the recovery rate for the central logical qubit in the $R=1/8$ Evenbly code is $44.02\pm 2.89\%$ in the $Y$ gauge. At first glance, this result may violate the quantum channel-capacity theorem [31], which sets a strict upper bound for erasure channel capacity at $43.75\%$ in our case. However, the channel capacity obtained is well within the error bounds in our simulation, and it is known that as long as physical error rates are below the threshold of the central logical qubit, all other bulk logical qubits can be decoded in parallel in holographic codes [22]. In this way, the central logical qubit’s recovery rate dictates the practical upper bound for all other bulk logical qubits. Our work therefore suggests that the word recovery rate for a constant-rate Evenbly code is only slightly lower than that of the central logical qubit, and may achieve channel capacity for finite-rate quantum erasure channels. Further work will be required in order to determine conclusively whether or not this is indeed the case.
@@
In addition to the code construction presented in the main text, we have also explored the threshold properties of other examples on $\{p,q\}$ hyperbolic tilings using the greedy recovery algorithm. These results are shown in Appendix C, and depict very interesting (if not similar) threshold data to the example fashioned in the main text. As stated in Appendices B and D, several other non-CSS seed-tensor constructions using AME and 2-uniform quantum states can be utilized to build an Evenbly code. These alternative Evenbly codes may be difficult to decode using the standard greedy decoder, owing to the lack of rotational invariance of the seed tensors. However, the Gaussian elimination or integer-optimization algorithms employed in this work could be employed; one may also consider modifying existing tensor-network decoders [22, 13] to accommodate Hadamard gates on the edges of the Evenbly code, as well as memory usage due to the exponentially-scaling number of sites observed in vertex inflation [44].
@@
As a final comment in this section, no prior work to our knowledge has yet considered fault-tolerance protocols [3] for a holographic code. Ref. [74] discovered the existence of a threshold in continuum AdS/CFT models, arising as a confinement/deconfinement phase transition. As the majority of holographic codes are stabilizer codes, one can minimally expect a generic holographic code to benefit from the implementation of a fault-tolerance protocol such as the flag-style protocols of [75, 76]. However, more elegant solutions may be possible; although not exhibiting bounded-weight check operators like in quantum low-density parity-check (qLDPC) codes [7], one may expect holographic codes to be amenable to similar fault-tolerance techniques, as in recent work demonstrating promising space and time overhead savings in the context of generalized concatenated codes [77, 78]. As providing for fault-tolerant operation of the code is a vital step towards the experimental realm, this topic is the subject of active research.
IV.3 Logical Gate Sets
@@
In this article, we have generalized the HTN ansatz to qubit-level codes, which we dub Evenbly codes. We have shown that this new, infinitely large code class exhibits unique and rich error correction properties that we have analyzed and reported on. Some of the highlights of this code include high erasure, depolarizing, and pure Pauli noise thresholds; low-weight transversal logical operators in the $X$ gauge; excellent distance scaling for bulk logical qubits; and constant-rate variants which are easy to tune and contrive. Lastly, our construction utilizes seed tensors which have already been prepared in the laboratory, implying that this class of holographic code may be amenable for experimental realization.
@@ Acknowledgements
We thank Aritra Sarkar, Charles Cao, and Jens Eisert for useful discussions. MS and SF thank the Intel corporation for financial support. AJ received funding from the Einstein Research Unit “Perspectives of a quantum digital transformation”. RH was supported by the Australian Research Council Centre of Excellence for Engineered Quantum Systems (Grant No. CE 170100009).
@@
MS and AJ developed the theoretical constructions and formalism for Evenbly codes. AJ developed and implemented the greedy reconstruction algorithm. JF and MS constructed the Gaussian elimination algorithm. JF constructed and implemented the distance-scaling approximations, as well as the integer-optimization decoder; this work was conducted during the master thesis of JF, and was supervised by MS and SF. RH provided guidance on the construction of both the Gaussian elimination and the integer-optimization decoders. MS and AJ wrote the manuscript. DE, SF, and AJ supervised and coordinated the project, as well as providing insight and guidance during the writing process.
@@ Appendix A Multipartite Maximally-Entangled States & Quantum Codes
MME states are highly-entangled many-body states in which certain reductions of the state are maximally mixed [19]; such states have found numerous applications in the construction of quantum error correction codes and quantum-communications protocols, among others [29, 20, 64]. In the case of systems consisting of two qubits, the Bell state |$\ket{\Psi^{+}_{2}}=\frac{1}{\sqrt{2}}(\ket{00}+\ket{11})$ is well-known as the simplest quantum state which is maximally entangled, since the state obtained after performing a partial trace on either of the two subsystems results in a maximally-mixed quantum state $\text{tr}_{A}\big{[}\ket{\Psi^{+}_{2}}\bra{\Psi^{+}_{2}}\big{]}=\mathbb{I}_{2}/2$. This state is locally equivalent to other states of the form $(U_{A}\otimes U_{B})\ket{\Psi_{2}^{+}}$, where $U_{A},U_{B}\in U(2)$. In a similar way, one can define a generalized Bell state for any bipartite quantum system with $n$ levels:
|$\ket{\Psi^{+}_{n}}=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\ket{i}_{A}\otimes\ket{i}_{B}.$ (27)
Here, one can easily check that the resulting RDMs for subsystem $A$ or $B$ will indeed be maximally mixed, as the entanglement entropy $\mathcal{S}$ of such states can be expressed as $\mathcal{S}(\ket{\Psi^{+}_{n}}\ket{\Psi^{+}_{n}})=\log{n}$ ).
@@
In vein with the notion of bipartite maximally-mixed reductions, $k$-uniform states are defined as pure states of $n$ distinguishable qudits in which all k-qudit reductions of the entire system are maximally mixed [85, 48, 49, 86, 50], with $k\leq\lfloor\frac{n}{2}\rfloor$. It is well-known that $k$-uniform states themselves are dual descriptions of stabilizer codes [49, 86], and that several inequivalent methods for constructing $k$-uniform states exist [48, 86, 49]. More specifically, a $k$-uniform state |$\ket{\psi_{\text{$k$-uniform}}}$ in a Hilbert space $\mathcal{H}(n,q)\in\mathbb{C}^{\otimes n}_{\chi}$ (where $\chi$ represents the local Hilbert-space dimension of all $n$ qudits), denoted by $k$-uniform(n,q) exists when the following condition holds:
$\text{tr}_{\mathcal{\bar{A}}}\big{[}\ket{\psi_{\text{{\hbox{k}}-uniform}}}\bra{\psi_{\text{{\hbox{k}}-uniform}}}\big{]}\propto\mathbb{I},$ (28)
@@
    Commutation: every pair of elements $\mathcal{G}_{i},\mathcal{G}_{j}$ must commute, i.e. $[\mathcal{G}_{i},\mathcal{G}_{j}]=0,\forall i,j\in\{1,...,m\}$, and
@@
    $k$-uniformity: the product $\mathcal{G}^{j_{1}}_{1}\cdots\mathcal{G}^{j_{m}}_{m}$ must result in an $n$-qubit Pauli operator containing no more than $(n-k-1)$ Pauli $I$ operators.
@@
It is additionally well-known that pure $k$-uniform states themselves form a set of quantum codes, formed as superpositions of classical maximum distance-separable (MDS) codes [49, 86]. A particular class of quantum codes known as stabilizer quantum codes can be modified via a technique known as shortening, and one can show the existence of a pure quantum stabilizer code whose spanning vectors are $(k-1)$-UNI states. Such codes can be defined for any local-dimension Hilbert space, and correspond to qudit-based stabilizer codes.
@@
The stabilizer $\mathcal{S}$ of a code is an Abelian subgroup of the $n$-fold Pauli group $\prod^{n}$, and does not contain the operator $I^{\otimes n}$; the simultaneous $+1$ eigenspace of operators forms what is known as the codespace. A stabilizer code exhibits a minimal representation of its stabilizer $\mathcal{S}$ in terms of $(n-k)$ independent generators:
@@ Appendix B An Evenbly code from a Graph State
Figure 10: A $\{5,4\}$ HTN qubit-level code. A) Complying with the constraints requires that any pair of neighboring tensors $A^{\prime}$ with an intermediate tensor $B$ must contract to unitaries $U^{\prime}_{1}$ and $U^{\prime}_{2}$ along two directions. B)-C) As described in Equation 35, the tensor $A^{\prime}$ is built from the amplitudes of the 5-qubit graph state, shortened into a $\llbracket 4,1,2\rrbracket$ code; in this graph-theoretic form, vertices (in green) denote qubits initialized as $\ket{+}$, and edges represent two-qubit CZ gates. As it is known that this quantum code is not invariant under planar index rotations, the tensors $A^{\prime}$ have an inherent directionality, denoted by a black arrow. D) Depending on the relative orientations of neighboring $A^{\prime}-B-A^{\prime}$ tensors concurrent in a given layer, one of three $B$ tensors is needed to fulfill the conditions in A), here shown up a normalization $\frac{1}{\sqrt{2}}$.
For convenience, we choose the physical state encoding the |+¯=|$\ket{\bar{+}}=\frac{\ket{\bar{0}}+\ket{\bar{1}}}{\sqrt{2}}$ logical state, which happens to coincide with the (AME) 5-qubit cycle graph state [89, 90]. This logical state is the unique ground state of the stabilizer Hamiltonian
@@ Appendix C Generalizations of Evenbly codes for Other Hyperbolic Tilings
Figure 11: Greedy recovery probabilities $p_{L}$ for Evenbly codes on more general $\{p,q\}$ tilings under erasure noise with probability $p_{\text{erasure}}$. We only consider the zero-rate code with $L$ layers of vertex inflation, with a single logical qubit, under various gauges. (a)-(b): The $q=4$ code considered in the main text can be placed on $\{p,4\}$ tilings for any $p\geq 5$. In the $Y$ gauge, recovery under the greedy algorithm does not appear to depend strongly on $p$, though in the $Z$ gauge the threshold seems to slightly increase with $q$. (c)-(d) A similar analysis can be performed for the $q=6$ code for any $p\geq 4$, for which greedy reconstruction under the $X$ gauge becomes non-trivial and equivalent to the $Y$ gauge. Similar to the $q=4$ code, increasing $p$ does not significantly alter $p_{L}$ at fixed $L$ under $X$/$Y$ gauge but improves it under $Z$ gauge. For these graphs, we have not shown the zeroth layer ($L=0$) with a single seed tensor, as it exhibits the same recovery curve for any setting (cf. Figure 7).
@@ Appendix D Alternative Evenbly code Constructions
Here, we present several constructions of alternative HTN codes, using non-CSS constructions of vertex tensors $A^{\prime}$, as well as the typical $B=H$ edge tensor condition. In addition to the two constructions provided in Sections III.2 and B, one may also formulate such alternative HTN quantum codes using the complex amplitudes from full 2-uniform states as the $A^{\prime}$ tensors. Previous work has shown that it is possible to generate new stabilizer quantum error correction codes from old ones by utilizing a procedure known as code shortening [2, 49, 88]. Here, we show that by employing $(q+1)$-partite 2-uniform states, one can perform logical implantation in the bulk using such 2-uniform quantum states. The implanted 2-uniform state can be then “shortened” from a $\llbracket q+1,0,k+1\rrbracket$ code to a $\llbracket q,1,k+1\rrbracket$ stabilizer code. We provide an example originating from a $\{5,5\}$ tiling below.
@@
$\mathcal{G}_{1}$ $=XXYYZZ,$ (37)
$\mathcal{G}_{2}$ $=XXZZYY,$ (38)
$\mathcal{G}_{3}$ $=XZZXXZ,$ (39)
$\mathcal{G}_{4}$ $=XYYXIZ,$ (40)
$\mathcal{G}_{5}$ $=YXIZXY,$ (41)
$\mathcal{G}_{6}$ $=YYYYII,$ (42)
@@
$\bar{Z}_{\llbracket 6,0,3\rrbracket}$ $=XXYYZZ=\mathcal{G}_{1},$ (43)
$\bar{X}_{\llbracket 6,0,3\rrbracket}$ $=XXZZYY=\mathcal{G}_{2}.$ (44)
The new stabilizer generators $\{\bar{\mathcal{G}}_{l}\}$, where $l=(n-k)-2=4$, are defined as
$\bar{\mathcal{G}}_{1}$ $=\mathcal{G}_{3}\circ\bar{Z}_{\llbracket 6,0,3\rrbracket}=IYXZYI,$ (45)
$\bar{\mathcal{G}}_{2}$ $=\mathcal{G}_{4}\circ\bar{Z}_{\llbracket 6,0,3\rrbracket}=IZIZZI,$ (46)
$\bar{\mathcal{G}}_{3}$ $=\mathcal{G}_{5}\circ\bar{X}_{\llbracket 6,0,3\rrbracket}=ZIZIZI,$ (47)
$\bar{\mathcal{G}}_{4}$ $=\mathcal{G}_{6}=YYYYII.$ (48)
Our new set of generators is now formed simply by removing the last (identity) Pauli from the full set described above, and are given by the stabilizer $\mathcal{S}_{\llbracket 5,1,2\rrbracket}=\{IYXZY,IZIZZ,ZIZIZ,YYYYI\}$; consequently, the new logical operators are $\bar{X}_{\llbracket 5,1,2\rrbracket}=XXZZY$, and $\bar{Z}_{\llbracket 5,1,2\rrbracket}=XXYYZ$. As in Section III.2, one may apply Hadamard gates to each of the boundary qubits for the stabilizers and logical operators, effectively flipping $X\mapsto Z$, and $Z\mapsto X$, as expected. We also checked numerically that the isometry constraints of Definition 2 are upheld, with the only difference being that rotational invariance is not upheld for this 2-uniform state.

### v2

@@ 4.2 Error Correction
The results conveyed demonstrate that Evenbly codes achieve similar logical recovery rates when compared to both topological and quantum low-density parity-check (qLDPC) codes [7, 6] in the code-capacity regime, as well as other holographic codes [13, 22, 28, 9, 8]. The rate of an Evenbly code is also tunable, a property generally ascribed to holographic codes [8, 14]. However, in contrast to the surface or color codes that can achieve a distance scaling $d=O(\sqrt{n})$ for a single logical qubit [41], the $\{5,4\}$ Evenbly code exhibits an $O(n^{0.67})$ scaling. This distance scaling can be subtle for Evenbly codes with multiple logical qubits, however: Though the weight of logical operators of a single logical bulk qubit (i.e., its bit distance $d_{\text{bit}}$) still generally scales as $O(n^{0.67})$ for a $\{5,4\}$ code, it is highly dependent on the bulk location and can be as low as $d_{\text{bit}}=2$. Giving more precise details on what types of lower and upper bounds can be expected for given orientation patterns of bulk logical qubits in Evenbly codes is in general still an open problem, and is the subject of ongoing research. Another important extension along this direction would be the examination of phenomenological noise models (which incorporate measurement errors); we leave this idea for future work.
@@
Additionally, we reported that the recovery rate for the central logical qubit in the $R=1/8$ Evenbly code is $44.02\pm 2.89\%$ in the $Y$ gauge. At first glance, this result may violate the quantum channel-capacity theorem [31], which sets a strict upper bound for erasure channel capacity at $43.75\%$ in our case. However, the channel capacity obtained is well within the error bounds in our simulation, and it is known that as long as physical error rates are below the threshold of the central logical qubit, all other bulk logical qubits can be decoded in parallel in holographic codes [22]. In this way, the central logical qubit’s recovery rate dictates the practical upper bound for all other bulk logical qubits. Our work therefore suggests that the word recovery rate for a constant-rate Evenbly code is only slightly lower than that of the central logical qubit, and may achieve channel capacity for finite-rate quantum erasure channels. In upcoming work, we determine conclusively whether or not this is indeed the case for the entire class of holographic quantum codes with finite rate [78].
@@
In addition to the code construction presented in the main text, we have also explored the threshold properties of other examples on $\{p,q\}$ hyperbolic tilings using the greedy recovery algorithm. These results are shown in Appendix C, and depict very interesting (if not similar) threshold data to the example fashioned in the main text. As stated in Appendices B and D, several other non-CSS seed-tensor constructions using AME and 2-uniform quantum states can be utilized to build an Evenbly code. These alternative Evenbly codes may be difficult to decode using the standard greedy decoder, owing to the lack of rotational invariance of the seed tensors. However, the Gaussian elimination or integer-optimization algorithms employed in this work could be used there; one may also consider modifying existing tensor-network decoders [22, 13] to accommodate Hadamard gates on the edges of the Evenbly code, as well as memory usage due to the exponentially-scaling number of sites observed in vertex inflation [46].
@@
As a final comment in this section, no prior work to our knowledge has yet considered fault-tolerant protocols [3] for a holographic code, although fault-tolerant preparation of precursor $k$-uniform states has recently been studied [79]. Ref. [80] discovered the existence of a threshold in continuum AdS/CFT models, arising as a confinement/deconfinement phase transition. As the majority of holographic codes are stabilizer codes, one can minimally expect a generic holographic code to benefit from the implementation of a fault-tolerant syndrome-extraction protocol such as the flag-style protocols of [81, 82]. However, more elegant solutions may be possible, such as those from [83]; although not exhibiting bounded-weight check operators like in quantum low-density parity-check (qLDPC) codes [7], one may expect holographic codes to be amenable to techniques similar to those from recent work demonstrating excellent space and time overhead savings in the context of generalized concatenated codes [84, 85, 86]. As providing for fault-tolerant operation of the code is a vital step towards the experimental realm, this topic is the subject of active research.
4.3 Logical Gate Sets
@@
In this article, we have generalized the HTN ansatz to qubit-level codes, which we dub Evenbly codes. We have shown that this new, infinitely large code class exhibits unique and rich error correction properties that we have analyzed and reported on. Some of the highlights of this code include: high erasure, depolarizing, and pure Pauli noise thresholds; low-weight transversal logical operators in the $X$ gauge; excellent distance scaling for bulk logical qubits; and constant-rate variants which are easy to tune and contrive. Lastly, our construction utilizes seed tensors which have already been prepared in the laboratory, implying that this class of holographic code may be amenable for experimental realization.
@@ Acknowledgements
We thank Aritra Sarkar, Charles Cao, and Jens Eisert for useful discussions. MS and SF thank the Intel corporation for financial support. AJ received funding from the Einstein Research Unit “Perspectives of a quantum digital transformation”. RH was supported by the Australian Research Council Centre of Excellence for Engineered Quantum Systems (Grant No. CE 170100009). DE was supported by the JST Moonshot R&D program under Grant JPMJMS226C.
@@
MS and AJ developed the theoretical constructions and formalism for Evenbly codes. AJ developed and implemented the greedy reconstruction algorithm. JF and MS constructed the Gaussian elimination algorithm. JF constructed and implemented the integer-optimization decoder during his master thesis, and was supervised by MS and SF. RH provided guidance for writing the Gaussian elimination and the integer-optimization decoders. MS and AJ wrote the manuscript. DE, SF, and AJ supervised and coordinated the project, as well as providing insight and guidance during the writing process.
@@ Appendix A Multipartite Maximally-Entangled States & Quantum Codes
[…]
@@
[…]
@@
    Commutation: every pair of elements $\mathcal{G}_{i},\mathcal{G}_{j}$ must commute, i.e. $\mathcal{G}_{i},\mathcal{G}_{j}]=0,\forall i,j\in\{1,...,m\}$, and
@@
    $k$-uniformity: the product $\mathcal{G}^{j_{1}}_{1}\cdots\mathcal{G}^{j_{m}}_{m}$ must result in an $n$-qubit Pauli operator containing no more than (nk1)(n-k-1) Pauli $I$ operators.
@@
It is additionally well-known that pure $k$-uniform states themselves form a set of quantum codes, formed as superpositions of classical maximum distance-separable (MDS) codes [51, 95]. A particular class of quantum codes known as stabilizer quantum codes can be modified via a technique known as shortening, and one can show the existence of a pure quantum stabilizer code whose spanning vectors are (k1)(k-1)-UNI states. Such codes can be defined for any local-dimension Hilbert space, and correspond to qudit-based stabilizer codes.
@@
The stabilizer $\mathcal{S}$ of a code is an Abelian subgroup of the $n$-fold Pauli group $\prod^{n}$, and does not contain the operator $I^{\otimes n}$; the simultaneous +1+1 eigenspace of operators forms what is known as the codespace. A stabilizer code exhibits a minimal representation of its stabilizer $\mathcal{S}$ in terms of (nk)(n-k) independent generators:
@@ Appendix B An Evenbly code from a Graph State
[…]
@@
Figure 11: Greedy recovery probabilities $p_{L}$ for Evenbly codes on more general $\{p,q\}$ tilings under erasure noise with probability $p_{\text{erasure}}$. We only consider the zero-rate code with $L$ layers of vertex inflation, with a single logical qubit, under various gauges. (a)-(b): The $q=4$ code considered in the main text can be placed on $\{p,4\}$ tilings for any $p\geq 5$. In the $Y$ gauge, recovery under the greedy algorithm does not appear to depend strongly on $p$, though in the $Z$ gauge the threshold seems to slightly increase with $q$. (c)-(d) A similar analysis can be performed for the $q=6$ code for any $p\geq 4$, for which greedy reconstruction under the $X$ gauge becomes non-trivial and equivalent to the $Y$ gauge. Similar to the $q=4$ code, increasing $p$ does not significantly alter $p_{L}$ at fixed $L$ under XX/$Y$ gauge but improves it under $Z$ gauge. For these graphs, we have not shown the zeroth layer ($L=0$) with a single seed tensor, as it exhibits the same recovery curve for any setting (cf. Figure 7).
@@ Appendix D Alternative Evenbly code Constructions
Here, we present several constructions of alternative Evenbly codes, using non-CSS constructions of vertex tensors $A^{\prime}$, as well as the typical $B=H$ edge tensor condition. In addition to the two constructions provided in Sections 3.2 and B, one may also formulate such alternative Evenbly quantum codes using the complex amplitudes from full 2-uniform states as the $A^{\prime}$ tensors. Previous work has shown that it is possible to generate new stabilizer quantum error correction codes from old ones by utilizing a procedure known as code shortening [2, 51, 97]. Here, we show that by employing $(q+1)$-partite 2-uniform states, one can perform logical implantation in the bulk using such 2-uniform quantum states. The implanted 2-uniform state can be then “shortened” from a $\llbracket q+1,0,k+1\rrbracket$ code to a $\llbracket q,1,k+1\rrbracket$ stabilizer code. We provide an example originating from a $\{5,5\}$ tiling below.
@@
$\mathcal{G}_{1}$ $=XXYYZZ,$ (35a)
$\mathcal{G}_{2}$ $=XXZZYY,$ (35b)
$\mathcal{G}_{3}$ $=XZZXXZ,$ (35c)
$\mathcal{G}_{4}$ $=XYYXIZ,$ (35d)
$\mathcal{G}_{5}$ $=YXIZXY,$ (35e)
$\mathcal{G}_{6}$ $=YYYYII,$ (35f)
@@
$\bar{Z}_{\llbracket 6,0,3\rrbracket}$ $=XXYYZZ=\mathcal{G}_{1},$ (36a)
$\bar{X}_{\llbracket 6,0,3\rrbracket}$ $=XXZZYY=\mathcal{G}_{2}.$ (36b)
The new stabilizer generators $\{\bar{\mathcal{G}}_{l}\}$, where l=(nk)2=4l=(n-k)-2=4, are defined as
$\bar{\mathcal{G}}_{1}$ $=\mathcal{G}_{3}\circ\bar{Z}_{\llbracket 6,0,3\rrbracket}=IYXZYI,$ (37a)
$\bar{\mathcal{G}}_{2}$ $=\mathcal{G}_{4}\circ\bar{Z}_{\llbracket 6,0,3\rrbracket}=IZIZZI,$ (37b)
$\bar{\mathcal{G}}_{3}$ $=\mathcal{G}_{5}\circ\bar{X}_{\llbracket 6,0,3\rrbracket}=ZIZIZI,$ (37c)
$\bar{\mathcal{G}}_{4}$ $=\mathcal{G}_{6}=YYYYII.$ (37d)
Our new set of generators is now formed simply by removing the last (identity) Pauli from the full set described above, and are given by the stabilizer $\mathcal{S}_{\llbracket 5,1,2\rrbracket}=\{IYXZY,IZIZZ,ZIZIZ,YYYYI\}$; consequently, the new logical operators are $\bar{X}_{\llbracket 5,1,2\rrbracket}=XXZZY$, and $\bar{Z}_{\llbracket 5,1,2\rrbracket}=XXYYZ$. As in Section 3.2, one may apply Hadamard gates to each of the boundary qubits for the stabilizers and logical operators, effectively flipping $X\mapsto Z$, and $Z\mapsto X$, as expected. We also checked numerically that the isometry constraints of Definition 2 are upheld, with the only difference being that rotational invariance is not upheld for this 2-uniform state.